OpenStudy (anonymous):

s=2πr^2+2πrh solve for h step by step please =]

OpenStudy (anonymous):

We have: \[ s=2\pi r^2+2\pi rh \]We combine like terms to find: \[ s=2\pi r(r+h) \implies \frac{s}{2\pi r}=r+h \implies \frac{s}{2\pi r}-r=h \]And we are done.